27 April 2001
N.B. Palmer

Hello Jose et al.

We are now at -64 13.010° S; -66 15.938° W (23:09 local) and steaming on a southeasterly course (133) at about 9 kts headed for Palmer Station. The wind is blowing around 30 kts and the temperature is about 0° C.

Today was a bit of one step forward and two steps back for the BIOMAPER-II group. We have been working to track down the noise problems in the BIOMAPER-II video system and Scott Gallager was down in the van on the deck just before lunch doing some testing on the fish when a large wave came onto the stern deck area and inundated the winch and washed up against the van. Immediately after the power to BIOMAPER-II failed because water got into a junction box

mounted on the side of the winch and shorted out a number of power wires i.e., the terminal strips to which the wires were screwed got fried along with some of the wires. The seas were big today because we have had 30 to 40 kt winds almost continuously since late last night. This made working on deck impossible without slowing the vessel down and heading into the seas. Also the air temperature was well below freezing. So after lunch, a number of us suited up in our boots and mustang suits and once on the deck determined in a few minutes that the smaller

junction box was the one with the problem. It took more than 2 hours to replace the terminal strips and reattach all of the wires. Scott Gallager took the lead in making all of the connections and rewiring the box with assistance principally from Mark Dennett, and Andy Girard. Matt Burke was the Marine Tech that provided assistance and stood watch as work ensued. Although the ship was moving slowly through the rough seas, occasional waves would slap up against the starboard rail and a several inches of water would come rushing across the deck to where the repair was taking place. Around three this afternoon we were finally able to power the system up and to our relief it worked.

During the day, the Hofmann group continued to drop XBT's at 10 nm intervals as we steamed across the southern portion of the Drake Passage. Several interesting features appeared in the temperature section and in the concurrently measured current field as determined from the ADCP (Acoustic Doppler Current Profiler). After we crossed the north wall of the polar front yesterday, we entered a region with a layer of cold water (around -1.0° C) ranging between 75 to 150 meters and centered about 100 m. This was left over "winter water" from last year's winter mixed layer. Surface currents (speeds of ~ 40 to 50 cm/sec) in this region were very strong to the east. The flow field abruptly reversed (flow to the west at speeds of ~30 to 40 cm/sec) at ~60 30'S as we cut across what appeared to be a meander of the polar front (or a large cyclonic

warm-core ring or eddy) . Further along at about 61° S, we again entered the polar frontal zone, a region where the winter water layer was present and the currents were again to the east albeit reduced in speed (~5 to 10 cm/sec).

This evening as we approached the Western Peninsula's shelf edge and the passive listening mooring SIOS2, we crossed the Southern Antarctic Circumpolar Current Front (~64° S). The front is defined as the point where the subsurface layer of 2° C water disappears and is replaced by colder water. We also crossed the Southern ACC boundary at 64 5'S, before moving up onto the shelf. This boundary is defined by the disappearance of 1.8° C water and replacement by colder water at the same depths. This crossing also marked the end point of this XBT transect.

As we passed over mooring SIOS2, Catherine Berchock dropped a Sonabouy (#2) in the water to calibrate the passive listening instrumentation sitting on the bottom (~3000 m depth). If whale calls were heard by both the bottom mounted instrument and the Sonabuoy, this would be very helpful in interpreting the mooring records. Even without marine mammal calls, the ship's noisy passage, which is recorded by both instruments, will assist in the calibration. During the passage over the mooring site, a CTD was dropped to provide information that can be used to compute a sound velocity profile. An attempt was made to drop an XCTD, but this was a first try and the software failed to record any data from the first drop and a second try was abandoned after it was clear that there was a problem with the way the software was setup which needed more time to solve.

Bird observations by Erik Chapman and Chris Ribic included sightings of many Cape petrels and Antarctic petrels, a few Blue petrels, and one Snow Petrel, and a moderate number of Southern fulmars (sometimes called the Antarctic fulmar). These beautiful birds were soaring along side the ship and appearing to take advantage of the 30 to 40 kt winds and the updrafts and other wind variations associated with the ship. There were no sightings of marine mammals today.

Tomorrow in early morning daylight, we plan to rendezvous with the Gould at Palmer Station to transfer gear needed by the Gould for their process work and some other material. Later in the day, both ships should be able to set off towards the first survey station to the southwest of the Palmer Station and the first process work site.

Cheers, Peter